Hi Can someone please help me this one I have been stuck on it for ages
0.01 = 0.995842 * 0.1180 / SQRTx

Answer by mananth(16949)

0.01 = 0.995842 * 0.1180 / SQRTx
multiply by sqrtx on both sides
sqrtx*0.01 = 0.005842*0.1180
sqrtx = 0.005842*0.1180/0.01
square both sides
x= (0.005842*0.1180/0.01)^2
x= 0.004752